Majer Corporation makes a product with the following standard costs:

Standard Quantity
or Hours
Standard Price or
Rate
Standard Cost Per Unit
Direct materials 6.8 ounces $ 3.00 per ounce $ 20.40
Direct labor 0.6 hours $ 15.00 per hour $ 9.00
Variable overhead 0.6 hours $ 3.00 per hour $ 1.80

The company reported the following results concerning this product in February.

Originally budgeted output 5,400 units
Actual output 5,000 units
Raw materials used in production 30,500 ounces
Actual direct labor-hours 7,930 hours
Purchases of raw materials 32,900 ounces
Actual price of raw materials $ 62.90 per ounce
Actual direct labor rate $ 72.40 per hour
Actual variable overhead rate $ 4.50 per hour

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The materials quantity variance for February is:
Answer #1

Materials Quantity Variance = (SQ-AQ) * SP

= (5000*6.8 - 30,500) * 3

= 3500 * 3

= 10,500 FAVOURABLE

Comment if you face any